# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v3

@@ III-C3 Numbers Task Metrics
The variation of width from one answer to the next. It is computed through a linear regression between the width of each answer and the question number. The metric itself represents the slope of this regression line. The variation in width is aimed to measure the presence of micrographia [42].
@@
The average answer pressure metric is a measure of the amount of pressure that a person applies while writing the answer. This metric can be leveraged as a time series metric to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The metric itself represents the slope of this regression line. This metric might help pick up dystonia metrics, involuntary muscle contractions commonly seen in neurodegenerative diseases.
@@
The total duration metric measures the time a participant spends answering one of the questions. Given the absence of a predefined time limit for the Numbers task, this metric provides insight into the participant's pace in completing the task.
III-C4 Writing Tasks Metrics
In this section, we present the metrics extracted from the writing tasks: freewrite, copy text, and copy text memory[49]. The computation of these metrics is shown in Figure 7.
Figure 7: Schematic of the metrics used in writing tasks, illustrated for the words 'This is an example'.
@@ III-C4 Writing Tasks Metrics
The pressure standard deviation metric is a measure of the spread of the amount of pressure a person applies while writing. This measures how steady the hand is when writing by capturing the variations in pressure. Only the nonzero pressure values were considered when calculating this metric, as no pressure can be computed if the pen is not touching the tablet.
@@
The average time between writing measures how long it takes for the participant to start writing after finishing a word. This shows how fluent the writing is and how stable the motor planning activity is. The metric was calculated by averaging the total time the pen was not touching the tablet.
@@
This metric was calculated by finding the difference between the $X$ value of when the pen is lifted and when it is put back down. Only the $X$ distance was considered because the $Y$ distance between the words is dependent on both the letters they are writing and how they write the letters.
@@
The total time metric gauges the duration required for task completion. It is determined by calculating the time elapsed from when the participant initially places the pen to start writing until they lift the pen after writing the last word. This metric holds greater significance for tasks like Copy Text and Copy Read Text, where participants' required text to write is the same.
@@ III-D Statistical Analysis
After extracting a total of 76 metrics (i.e., 54 task-agnostic metrics and 22 task-specific metrics), we aim to evaluate their relevance for assessing and characterizing neurodegenerative diseases (NDs). Our objective is twofold: first, to identify which metrics effectively differentiate the CTL group from each of the disease groups (AD, PD, and PDM), and second, to assess whether handwriting metrics can distinguish PD from other forms of Parkinsonism (PDM).
To this end, we perform a Kruskal-Wallis H-test across the four experimental groups (CTL, AD, PD, PDM) for each metric, using the scipy.stats.kruskal function from the Scipy toolkit [^2] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kruskal.html. For metrics showing a significant global effect (p<<0.05), we apply Dunn’s post hoc test to identify specific group differences. To control for false discoveries arising from multiple testing, we apply False Discovery Rate (FDR) correction using the statsmodels.stats.multitest.fdrcorrection function [^3] https://www.statsmodels.org/dev/generated/statsmodels.stats.multitest.fdrcorrection.html across the set of p-values grouped by task and test. Since we test 14 tasks across 4 pairwise group comparisons, we apply $4\times 14=56$ FDR corrections. To complete our analysis of the separability of those groups of participants, we measured the Area Under the Curve (AUC) for pairs of participant groups, which will be presented alongside FDR-corrected p-values.
Additionally, to investigate how individual metrics relate to cognitive and motor functioning, we compute the Pearson correlation coefficient between each of the 76 extracted features and both the MoCA [34] (Montreal Cognitive Assessment, for all participants) and UPDRS-III [35] (Unified Parkinson's Disease Rating Scale, motor section, only measured for the PD and PDM participants) scores across participants. To provide a more detailed analysis of motor function, we also compute correlations between each feature and the individual subdomains of the UPDRS-III score, including: rigidity, upper extremity bradykinesia, lower extremity bradykinesia, arising from a chair, gait, freezing of gait, postural stability, posture, gait and posture composite, kinetic tremor, postural tremor, resting tremor, global tremor score, and non-tremor components.
@@
III-D1 Grouped metrics
We hypothesized that certain metrics extracted for a specific type of task would exhibit consistent behavior across all tasks. To test this hypothesis, we grouped the tasks together. For each group, we conducted the same statistical comparisons outlined in section III-D. However, instead of analyzing the recordings from individual tasks, we considered all recordings within each task group. We then measured the significance of all metrics for each of these task groups as if they were a single task.
@@ III-D1 Grouped metrics
Table II summarizes the tasks at hand sorted per group and the specific metrics proposed per group of tasks.
TABLE II: summary of the tasks at hand and their specific metrics.
Group Tasks Specific Metrics
@@ III-E Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using metrics extracted from each task. Namely, it involves a binary classification (e.g., AD or CTL) that facilitates the assignment of samples to their corresponding groups, thereby enabling the evaluation of the metrics' effectiveness in distinguishing between the categories under investigation. As for the statistical analysis, the feature set used in the classification models consists of both task-specific and general metrics, ensuring a comprehensive representation of handwriting dynamics. The number of metrics varies across tasks, with Copycube, Drawclock, Copymage Memory, and Copymage containing 54 metrics, while Copyreadtext, Copytext, and Freewrite include 62 metrics. The Numbers task has 68 metrics, whereas Point DOM and Point NONDOM each contain 28 metrics, and Point Sustained has 27 metrics. Additionally, Spiral DOM includes 38 metrics, while Spiral NONDOM and Spiral Pataka each have 37 metrics. These metrics capture various aspects of handwriting, including kinematics (e.g., velocity, acceleration, jerk), spatial properties (e.g., trajectory deviations, area), and pressure dynamics (e.g., mean pressure, pressure variability).
@@
These experiments further evaluate the discriminative power of the metrics and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Random Forest (RF), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^4] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm. For each classifier, hyperparameter tuning was performed using nested cross-validation (NCV) to ensure optimal generalization performance. Specifically, for the BG classifier, we tuned max_samples (proportion of training samples used per estimator) and n_estimators (number of base estimators). For the MLP classifier, we optimized hidden_layer_sizes (number of neurons in the hidden layer). Finally, for the RF classifier, we tuned n_estimators (number of trees in the forest) and max_metrics (number of metrics considered per split). NCV is a widely used approach for assessing model performance, especially when dealing with limited data, as it helps provide an unbiased estimate of how well the model will generalize to new data. By incorporating two levels of data partitioning, NCV enables both model evaluation and hyperparameter tuning while minimizing overfitting and overly optimistic performance estimates [50, 51]. The NCV framework consists of two levels of cross-validation to ensure robust model evaluation and mitigate overfitting. In our implementation, the outer loop employs a 10-fold cross-validation, partitioning the dataset into 10 folds, with one fold held out for testing while the remaining nine are used for training. Within each outer training set, an inner 10-fold cross-validation is conducted to optimize hyperparameters. The best hyperparameter configuration is selected based on the average F1-score computed across all 10 inner folds, which helps mitigate variance in performance estimates and ensures stability in hyperparameter selection. The F1-score was chosen as the optimization criterion due to class imbalance. To prevent data leakage, cross-validation was strictly performed at the subject level. Speaker-independent folds were created to ensure that all samples from a given individual appeared only in either the training or test set within any iteration. This setup prevents data from the same subject from being in both training and testing, thereby avoiding overly optimistic performance estimates. The entire NCV process involved 100 hyperparameter tuning runs (10 outer folds $\times$ 10 inner folds), with final performance estimates derived by averaging results across the outer folds. This methodology provides a robust assessment of model generalization while maintaining strict speaker independence.
@@ IV Results
In this section, we first present the most significant metrics obtained by groups of metrics for each test realized. Then, we comment on the results obtained in the classification experiments for each task.
IV-A Task-agnostic Metrics Results
Out of 54 task-agnostic metrics computed, we found 53 were significant ($p\leq 0.05$) for, at least one task and one groups comparison. The complete table is available in the Appendix; however, we prioritize limited sets due to the considerable number of significant metrics. We present the metrics most effectively distinguishing AD and PD from control participants and separate PD from PDM participants. Subsequently, we will highlight metrics that exhibited robust behavior across multiple comparisons.
IV-A1 AD vs CTL
Over the 54 task-agnostic metrics, 50 are significant ($p\leq 0.05$) in comparing AD and CTL participants for at least one task, and 29 have p-values smaller than 0.001. We present the 10 metrics having the lowest p-values in Table III across all tasks. Our most significant and consistent metrics across all tasks are the Shannon entropy, the duration, the mean of the speed and accelerations for point tasks, and the variations of angular speed and acceleration for the drawing and writing tasks. We show the box plots for some of those metrics in Figure 9.
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one's handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear speed and linear acceleration, as well as duration, all show how significantly slower AD participants are from the control group, presumably due to processing inefficiency. Finally, the differences in angular speed and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant metrics were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD assessment, as opposed to the point and spiral tasks.
When comparing the tasks Copy Text and Copy Read Text, the difference in speed and horizontal acceleration is significantly higher for AD participants than CTL participants ($p$-value $\leq$ 0.001 for both), which shows how the multi-tasking has a way higher effect on cognitive participants.
@@ IV-A1 AD vs CTL
TABLE III: $p$-values and AUC for the best 10 metrics for AD vs CTL comparison. The table is repeated to fit the page. Significant $p$-values are bolded, as well as AUC values over 0.75
AD vs CTL Copy Image Copy Text Copy Read Text
@@
AD vs CTL Numbers Draw Clock Freewrite
@@
AD vs CTL Point Dominant Point Non Dominant Point Sustained
@@
in-air std speed 0.159 0.650 0.004 0.737 0.390 0.647
Figure 9: Boxplots of the most significant metrics (AD vs CTL) presented across drawing and writing tasks for all groups of participants.
@@ IV-A2 PD vs CTL
Over the 54 task-agnostic metrics, 10 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 7 have p-values under 0.001. Table IV shows the top 10 metrics with the lowest p-values across all tasks for the PD vs. CTL comparison.
TABLE IV: $p$-values and AUC for the best 10 metrics for PD vs CTL comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
@@
in-air avg. angular speed <<0.001 0.742 0.021 0.743 0.008 0.768
in-air avg. speed 0.031 0.608 1.000 0.620 0.437 0.630
in-air avg. X speed <<0.001 0.639 1.000 0.616 0.129 0.678
in-air avg. Y speed <<0.001 0.682 0.041 0.711 0.143 0.678
@@
in-air std angular speed <<0.001 0.733 0.021 0.756 0.009 0.759
These metrics primarily revolve around in-air dynamics, variations, and averages of speeds and angular speeds. They are notably discriminatory for point tasks. Additionally, we present box plots for these significant metrics in Figure 10 for further visualization and analysis.
Figure 10: Boxplots of some the best metrics extracted for PD vs CTL, all from angular speed and accelerations, presented for drawing and point tasks for all groups of participants.
For detecting PD, the most significant metrics were for drawing and point tasks. The point tasks were better suited to assess motoric metrics of PD, such as tremor and bradykinesia. Accordingly, we observed subjects with PD showed significantly lower angular speed and angular acceleration in air, which could be indicators of hypometria [52]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular speed is still a discriminative metric, but metrics of variation (as standard deviation or amplitudes) of the speed and acceleration on the tablet were notable. These findings may represent, tremor, dystonia, dyskinesia, myoclonus other other complicating motor symptoms that develop in some persons with PD [10]. However, none of the task-agnostic metrics measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.

### v4

@@ III-C3 Numbers Task Features
The variation of width from one answer to the next. It is computed through a linear regression between the width of each answer and the question number. The feature itself represents the slope of this regression line. The variation in width is aimed to measure the presence of micrographia [43].
@@
The average answer pressure feature is a measure of the amount of pressure that a person applies while writing the answer. This feature can be leveraged as a time series feature to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The feature itself represents the slope of this regression line. This feature might help pick up dystonia features, involuntary muscle contractions commonly seen in neurodegenerative diseases.
@@
The total duration feature measures the time a participant spends answering one of the questions. Given the absence of a predefined time limit for the Numbers task, this feature provides insight into the participant's pace in completing the task.
III-C4 Writing Tasks Features
In this section, we present the features extracted from the writing tasks: freewrite, copy text, and copy text memory[50]. The computation of these features is shown in Figure 7.
Figure 7: Schematic of the features used in writing tasks, illustrated for the words 'This is an example'.
@@ III-C4 Writing Tasks Features
The pressure standard deviation feature is a measure of the spread of the amount of pressure a person applies while writing. This measures how steady the hand is when writing by capturing the variations in pressure. Only the nonzero pressure values were considered when calculating this feature, as no pressure can be computed if the pen is not touching the tablet.
@@
The average time between writing measures how long it takes for the participant to start writing after finishing a word. This shows how fluent the writing is and how stable the motor planning activity is. The feature was calculated by averaging the total time the pen was not touching the tablet.
@@
This feature was calculated by finding the difference between the $X$ value of when the pen is lifted and when it is put back down. Only the $X$ distance was considered because the $Y$ distance between the words is dependent on both the letters they are writing and how they write the letters.
@@
The total time feature gauges the duration required for task completion. It is determined by calculating the time elapsed from when the participant initially places the pen to start writing until they lift the pen after writing the last word. This feature holds greater significance for tasks like Copy Text and Copy Read Text, where participants' required text to write is the same.
@@ III-D Statistical Analysis
After extracting a total of 76 features (i.e., 54 task-agnostic features and 22 task-specific features), we aim to evaluate their relevance for assessing and characterizing neurodegenerative diseases (NDs). Our objective is twofold: first, to identify which features effectively differentiate the CTL group from each of the disease groups (AD*, PD, and PDM), and second, to assess whether handwriting features can distinguish PD from other forms of Parkinsonism (PDM).
To this end, we perform a Kruskal-Wallis H-test across the four experimental groups (CTL, AD, PD, PDM) for each feature, using the scipy.stats.kruskal function from the Scipy toolkit [^2] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kruskal.html. When a participant had multiple sessions, the average of all sessions for each feature is used, to insure the independence of all observations. For features showing a significant global effect (p<<0.05), we apply Dunn’s post hoc test to identify specific group differences. To control for false discoveries arising from multiple testing, we apply global False Discovery Rate (FDR) correction across all tasks and features for each comparison, using the statsmodels.stats.multitest.fdrcorrection function [^3] https://www.statsmodels.org/dev/generated/statsmodels.stats.multitest.fdrcorrection.html. To complete our analysis of the separability of those groups of participants, we measured the Area Under the Curve (AUC) for pairs of participant groups, which will be presented alongside FDR-corrected p-values.
Additionally, to investigate how individual features relate to cognitive and motor functioning, we compute the Pearson correlation coefficient between each of the 76 extracted features and both the MoCA [35] (Montreal Cognitive Assessment, for all participants) and UPDRS-III [36] (Unified Parkinson's Disease Rating Scale, motor section, only measured for the PD and PDM participants) scores across participants. To provide a more detailed analysis of motor function, we also compute correlations between each feature and the individual subdomains of the UPDRS-III score, including: rigidity, upper extremity bradykinesia, lower extremity bradykinesia, arising from a chair, gait, freezing of gait, postural stability, posture, gait and posture composite, kinetic tremor, postural tremor, resting tremor, global tremor score, and non-tremor components.
@@
III-D1 Grouped features
We hypothesized that certain features extracted for a specific type of task would exhibit consistent behavior across all tasks. To test this hypothesis, we grouped the tasks together. For each group, we conducted the same statistical comparisons outlined in section III-D. However, instead of analyzing the recordings from individual tasks, we considered all recordings within each task group. We then measured the significance of all features for each of these task groups as if they were a single task.
@@ III-D1 Grouped features
Table II summarizes the tasks at hand sorted per group and the specific features proposed per group of tasks.
TABLE II: summary of the tasks at hand and their specific features.
Group Tasks Specific Features
@@ III-E Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using features extracted from each task. Namely, it involves a binary classification (e.g., AD* vs CTL) that facilitates the assignment of samples to their corresponding groups, thereby enabling the evaluation of the features' effectiveness in distinguishing between the categories under investigation. As for the statistical analysis, the feature set used in the classification models consists of both task-specific and general features, ensuring a comprehensive representation of handwriting dynamics. The number of features varies across tasks, with Copycube, Drawclock, Copymage Memory, and Copymage containing 54 features, while Copyreadtext, Copytext, and Freewrite include 62 features. The Numbers task has 68 features, whereas Point DOM and Point NONDOM each contain 28 features, and Point Sustained has 27 features. Additionally, Spiral DOM includes 38 features, while Spiral NONDOM and Spiral Pataka each have 37 features. These features capture various aspects of handwriting, including dynamics (e.g., velocity, acceleration, jerk), spatial properties (e.g., trajectory deviations, area), and pressure dynamics (e.g., mean pressure, pressure variability).
@@
These experiments further evaluate the discriminative power of the features and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Random Forest (RF), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^4] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm. For each classifier, hyperparameter tuning was performed using nested cross-validation (NCV) to ensure optimal generalization performance. Specifically, for the BG classifier, we tuned max_samples (proportion of training samples used per estimator) and n_estimators (number of base estimators). For the MLP classifier, we optimized hidden_layer_sizes (number of neurons in the hidden layer). Finally, for the RF classifier, we tuned n_estimators (number of trees in the forest) and max features (number of features considered per split). NCV is a widely used approach for assessing model performance, especially when dealing with limited data, as it helps provide an unbiased estimate of how well the model will generalize to new data. By incorporating two levels of data partitioning, NCV enables both model evaluation and hyperparameter tuning while minimizing overfitting and overly optimistic performance estimates [51, 52]. The NCV framework consists of two levels of cross-validation to ensure robust model evaluation and mitigate overfitting. In our implementation, the outer loop employs a 10-fold cross-validation, partitioning the dataset into 10 folds, with one fold held out for testing while the remaining nine are used for training. Within each outer training set, an inner 10-fold cross-validation is conducted to optimize hyper-parameters. The best hyperparameter configuration is selected based on the average F1-score computed across all 10 inner folds, which helps mitigate variance in performance estimates and ensures stability in hyperparameter selection. The F1-score was chosen as the optimization criterion due to class imbalance. To prevent data leakage, cross-validation was strictly performed at the subject level. Speaker-independent folds were created to ensure that all samples from a given individual appeared only in either the training or test set within any iteration. This setup prevents data from the same subject from being in both training and testing, thereby avoiding overly optimistic performance estimates. The entire NCV process involved 100 hyperparameter tuning runs (10 outer folds $\times$ 10 inner folds), with final performance estimates derived by averaging results across the outer folds. This methodology provides a robust assessment of model generalization while maintaining strict speaker independence. For our chosen model, the Bagging Classifier, the following list of hyper-parameters was explored, and the best combination for each task was kept:
  • max_samplesmax\_samples: 0.05, 0.1, 0.2, 0.5
  • n_estimatorsn\_estimators: 10, 100, 1000
The full list of explored hyper-parameters for every explored model is provided in the associated github [^5] The full code allowing for the reproduction of all experiments will be released upon acceptance on github.
@@ IV Results
In this section, we first present the most significant features obtained by groups of features for each test realized. Then, we comment on the results obtained in the classification experiments for each task.
IV-A Task-agnostic Features Results
Out of 54 task-agnostic features computed, we found 53 were significant ($p\leq 0.05$) for, at least one task and one groups comparison. The complete table is available in the Appendix; however, we prioritize limited sets due to the considerable number of significant features. We present the features most effectively distinguishing AD* and PD from control participants and separate PD from PDM participants. Subsequently, we will highlight features that exhibited robust behavior across multiple comparisons.
IV-A1 AD* vs CTL
Over the 54 task-agnostic features, 50 are significant ($p\leq 0.05$) in comparing AD* and CTL participants for at least one task, and 29 have p-values smaller than 0.001. We present the 10 features having the lowest p-values in Table III across all tasks. Our most significant and consistent features across all tasks are the Shannon entropy, the duration, the mean of the velocity and accelerations for point tasks, and the variations of angular velocity and acceleration for the drawing and writing tasks. We show the box plots for some of those features in Figure 9.
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one's handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear velocity and linear acceleration, as well as duration, all show how significantly slower AD* participants are from the control group, presumably due to processing inefficiency. Finally, the differences in angular velocity and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant features were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD* assessment, as opposed to the point and spiral tasks.
When comparing the tasks Copy Text and Copy Read Text, the difference in velocity and horizontal acceleration is significantly higher for AD* participants than CTL participants ($p$-value $\leq$ 0.001 for both), which shows how the multi-tasking has a way higher effect on cognitive participants.
@@ IV-A1 AD* vs CTL
TABLE III: $p$-values and AUC for the best 10 features for AD* vs CTL comparison. The table is repeated to fit the page. Significant $p$-values are bolded, as well as AUC values over 0.75
AD* vs CTL Copy Image Copy Text Copy Read Text
@@
AD* vs CTL Numbers Draw Clock Freewrite
@@
AD* vs CTL Point Dominant Point Non Dominant Point Sustained
@@
in-air std velocity 0.159 0.650 0.004 0.737 0.390 0.647
Figure 9: Boxplots of the most significant features (AD* vs CTL) presented across drawing and writing tasks for all groups of participants.
@@ IV-A2 PD vs CTL
Over the 54 task-agnostic features, 10 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 7 have p-values under 0.001. Table IV shows the top 10 features with the lowest p-values across all tasks for the PD vs. CTL comparison.
TABLE IV: $p$-values and AUC for the best 10 features for PD vs CTL comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
@@
in-air avg. angular velocity <<0.001 0.742 0.021 0.743 0.008 0.768
in-air avg. velocity 0.031 0.608 1.000 0.620 0.437 0.630
in-air avg. X velocity <<0.001 0.639 1.000 0.616 0.129 0.678
in-air avg. Y velocity <<0.001 0.682 0.041 0.711 0.143 0.678
@@
in-air std angular velocity <<0.001 0.733 0.021 0.756 0.009 0.759
These features primarily revolve around in-air dynamics, variations, and averages of velocitys and angular velocitys. They are notably discriminatory for point tasks. Additionally, we present box plots for these significant features in Figure 10 for further visualization and analysis.
Figure 10: Boxplots of some the best features extracted for PD vs CTL, all from angular velocity and accelerations, presented for drawing and point tasks for all groups of participants.
For detecting PD, the most significant features were for drawing and point tasks. The point tasks were better suited to assess motoric features of PD, such as tremor and bradykinesia. Accordingly, we observed subjects with PD showed significantly lower angular velocity and angular acceleration in air, which could be indicators of hypometria [53]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular velocity is still a discriminative feature, but features of variation (as standard deviation or amplitudes) of the velocity and acceleration on the tablet were notable. These findings may represent, tremor, dystonia, dyskinesia, myoclonus other other complicating motor symptoms that develop in some persons with PD [10]. However, none of the task-agnostic features measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.